OpenStudy (winsteria):

Melanie’s monthly pay is $4349.76. She usually uses 3/4 of that amount to pay her household expenses, and 1/12 of that amount to pay for tuition. She deposits the rest into her savings account. This month, Melanie deposited twice the amount of money into her savings account than she normally deposits. How much did she deposit into her savings account this month? CAN SOMEONE PLEASE HELP ME ILL GIVE YOU A MEDAL AND A FAN IF YOU AT LEAST HELP ME PLEASE

OpenStudy (anonymous):

Do you know how to set up the equation?

OpenStudy (winsteria):

No D:

OpenStudy (anonymous):

Allright I will walk you through it. First we want to start with 4349.76 because that is the total she gets per month. Next we need to subtract 3/4 of her monthly pay and 1/12 of her pay too. Are you following?

OpenStudy (winsteria):

Yes what else???

OpenStudy (anonymous):

So Our equation is set up like this 4349.76 - (10/12*4349.76)= awnser We can simplify the fraction into 10/12 because Remember how to set up your equation? 3/4 + 1/12=10/12

OpenStudy (anonymous):

Otherwise our equation would be longer... Remember how to set up your equation? 4349.76 - ((3/4 * 4349.76) + (1/12*4349.76))

OpenStudy (winsteria):

So I tried to get the answer can you tell me if I did it right or wrong, my answer is 724.96 ???

OpenStudy (anonymous):

That is correct. Hopefully now you can use this question and the previous as a reference for the others to come.

OpenStudy (winsteria):

Thank you so much you helped me a lot thank you!!!

OpenStudy (anonymous):

You helped yourself.

OpenStudy (anonymous):

Don't forget the problem was Melanie deposited twice the amount of money into her savings account than she normally deposits. So $724.96 x2 = 1449.92